OpenStudy (anonymous):

slope 1/3 passing though (5,4)

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

slope 1/3 passing though (5,4)

OpenStudy (anonymous):

cause if you're looking for an equation in slope-intercept form it would be- y=1/3x+4

OpenStudy (zzr0ck3r):

no it wouldn't @Mhm120

OpenStudy (anonymous):

it depends on the question technically.

OpenStudy (zzr0ck3r):

no it doesnt

OpenStudy (anonymous):

graph a line with slope 1/3 passing though (5,4)

OpenStudy (zzr0ck3r):

y = mx+b is the equation of a line, m = slope and b = y intercept you have \[y=\frac{1}{3}x+b\] because they gave you slope, you now need to find b we will us the point they gave (5,4) subbing in x=5 and y = 4 we get \[4=\frac{1}{3}5+b\]solve for b

OpenStudy (anonymous):

lol ohhh okay nvm forget what I said then

OpenStudy (anonymous):

that's different.

OpenStudy (anonymous):

7/3